# Data-driven inventory management for new products: A warm-start and adjusted Dyna-$Q$ approach

### v2

@@ III-A Problem Formulation and Environment Construction
The state variable contains vital information on inventory levels. The state at period $t$ is shown as: $s_{t}=\{s^{1}_{t},\,s^{2}_{t},\,s^{3}_{t}\}$, where $s^{i}_{t}$ represents the inventory levels of products that have $i$ days shelf-life. For instance, $s^{3}_{t}$ denotes the inventory level of products with 3-day shelf-life, which are the fresh products the store just received on that day. The state $s^{2}_{t}$ and $s^{1}_{t}$ denote the inventory level of products with 2 days or 1-day shelf-life, respectively. At the end of period $t$, the products with $s^{1}_{t}$ will be disposed of from the inventory. Note that $s^{1}_{t},s^{2}_{t}$ and $s^{3}_{t}$ are all integers.
The bakery’s action $a_{t}$ denotes the order quantity of products at the end of period $t$. Suppose that the maximum possible inventory level $|\mathbb{S}|$ is predetermined and the quantity $a_{t}$ is an integer and has a proper upper bound depending on the maximal possible demand and inventory level.
In each period, the total cost $C_{t}$ is represented as
@@
For our problem, two state transition processes will occur: one at the beginning of the period and the other at the end associated with customer demand. The first state transition is elaborated in the formula as $s^{1}_{t}=s^{2}_{t-1}$, $s^{2}_{t}=s^{3}_{t-1}$, $s^{3}_{t}=a_{t-1}$, which summarize that the shelf-life of all products will decrease by 1 due to decaying. Then the $s_{t}$ will be used to compute the cost by equation (6). The second state transition occurs after the bakery interacts with the environment, returning daily demand information $d_{t}$:
@@
The goal of the inventory management problem is to find an optimal policy (optimal $a_{t}$ given state $s_{t}$ at each period), such that the long-term expected cost $\mathbb{E}[\sum_{t=1}^{\infty}\gamma^{t-1}C_{t}]$ is minimized.
@@ III-B Implementation and Result Discussions
We first apply the basic adjusted Dyna-$Q$ algorithm (without warm-start), with parameters $\alpha=0.3,\,\gamma=0.9,\epsilon_{0}=0.4,\,\epsilon_{\min}=0.1,\,\,N_{0}=100,\,N_{\min}=10,\,\partial_{\epsilon}=7500,\,\partial_{N}=5000$. For comparison, we construct the model $M(s,\,a)$ by two either BNNs or Multi-Layer Perceptrons (MLPs) (one for cost function and another for state transition probability). MLP is a neural network consisting of multiple layers of interconnected neurons. Its parameters are optimized by back-propagation, and the prediction error is minimized by gradient descent [19]. Our proposed algorithm is implemented across 500 training episodes, and the average cost per iteration has been recorded. The training processes are recorded in Fig.2, which shows the average cost occurs per iteration, where we observe a stable cost at the end of the training horizon. Thus, the trained optimal policies are also stable.
@@
Next, we implement the adjusted Dyna-$Q$ algorithm with warm-start via forecasting and simulation. We investigate the following two scenarios, training and testing, respectively. Here, the bakery manager is required to form an ordering policy at an early stage after the new product is released, without prior knowledge of its demand distribution. We choose the parameters $\alpha=0.1,\,\gamma=0.9\,\partial_{\epsilon}=1000,\,\partial_{N}=1000$ and $\sigma^{2}=5$.
Scenario 1 (Training): In this scenario, we set: $\epsilon_{0}=0.4,\,\epsilon_{\min}=0,\,N_{0}=100,\,N_{\min}=0$. We first implement the method in Section II.$B$ with $h=10$ and incorporate the warm-start information into the model of adjusted Dyna-$Q$. Then the training starts right at iteration $t=1$, when the new product, Boule 400g, has just been released and its demand is observed. The training horizons are set as one month i.e., $T=30$. The total cost of the training will be recorded, and its average value will be over 100 training episodes. We also record the variance of total cost over all episodes. Table II records the experiment results where the “Shortage percentage” records the proportion of days within a month when the unfulfilled demand occurs.
@@
Algorithm Warm- start Average total cost Shortage percentage Daily average inventory holding Variance in total cost
@@
Algorithm Warm- start Average total cost Shortage percentage Daily average inventory holding Variance in total cost
@@
From the experimental results of Scenario 1 recorded in Table II, we observe that the warm-start information does not uniformly improve the algorithms’ performance. Regarding total inventory cost generated in the first month, the three algorithms do not show a distinct difference in Scenario 1, with or without warm-start information. Note that the classic Dyna-$Q$ without warm-start information is the most effective, with the total cost during the training period at 61.86. However, incorporating warm-start information helps to stabilize all the algorithms through total cost variance reduction by at least $31.6\%$.
From the experimental results of Scenario 2 recorded in Table III, we observe that the adjusted Dyna-$Q$ algorithm with or without warm-start information performs better than all the other algorithms. In addition, incorporating warm-start information helps the adjusted Dyna-$Q$ algorithm perform the best on testing. Another observation is that the algorithms with warm-start information perform more stably with significantly lower variance in total cost on testing. Finally, we observe the phenomenon that the adjusted Dyna-$Q$ algorithm produces policies reducing the shortage percentage by holding more inventory on average.
Compared with the results in Scenario 1, we find that the variance of the total cost is much more significantly reduced in Scenario 2, which shows the strong out-of-sample performance of the proposed approach and its strong reliability in real-life applications. Moreover, the adjusted Dyna-$Q$ also indicates a more significant effect in reducing the total cost compared with Dyna-$Q$ due to the decaying exploration probability and planning steps that can eliminate the discrepancy between the model and the new product’s demand distribution.
@@ IV Conclusion
The paper presents an RL algorithm for inventory management of newly launched products, particularly in scenarios where historical demand information is limited or unavailable. The proposed adjusted Dyna-Q algorithm builds upon the classic Dyna-Q algorithm, aiming to strike a balance between model-based and model-free approaches during training. Additionally, a demand forecasting model is introduced, leveraging historical demand data from existing similar products to provide warm-start information for the adjusted Dyna-Q algorithm.
A case study using real-life data is conducted to validate the proposed approach. The adjusted Dyna-Q algorithm demonstrates significant benefits, including up to a 23.7% reduction in average daily cost compared to Q-learning and up to a 77.5% decrease in training time compared to classic Dyna-Q. Two scenarios are presented for generating and incorporating warm-start information into the algorithm for both training and testing. The results show that the adjusted Dyna-Q algorithm achieves the lowest total cost and relatively low shortage percentages compared to other algorithms. Algorithms that incorporate warm-start information exhibit greater stability and lower variance in total cost during testing.
For future research directions, several possibilities are suggested. First, further testing of the proposed approach in various inventory management contexts and industries could assess its versatility and scalability. Second, evaluating the adaptability of the approach to specific uncertainty patterns, such as seasonal demand fluctuations or sudden market changes, could provide insights into its robustness. Lastly, enhancing the demand prediction model to measure similarity between existing products and the new product could improve the incorporation of historical data from similar products with high similarity during training.

### v3

@@ III-A Problem Formulation and Environment Construction
The state variable contains vital information on inventory levels. The state at period $t$ is shown as: $s_{t}=\{s^{1}_{t},\,s^{2}_{t},\,s^{3}_{t}\}$, where $s^{i}_{t}$ represents the inventory levels of products that have $i$ days shelf-life. For instance, $s^{3}_{t}$ denotes the inventory level of products with a 3-day shelf-life, the fresh products the store just received that day. The state $s^{2}_{t}$ and $s^{1}_{t}$ denote the inventory level of products with 2 days or 1-day shelf-life, respectively. At the end of period $t$, the products with $s^{1}_{t}$ will be disposed of from the inventory. Note that $s^{1}_{t},s^{2}_{t}$ and $s^{3}_{t}$ are all integers.
The bakery’s action $a_{t}$ denotes the order quantity of products at the end of period $t$. Suppose that the maximum possible inventory level $|\mathbb{S}|$ is predetermined and the quantity $a_{t}$ is an integer with a proper upper bound depending on the maximum possible demand and inventory level. In each period, the total cost $C_{t}$ is represented as
@@
For our problem, two state transition processes will occur: one at the beginning of the period and the other at the end associated with customer demand. The first state transition is elaborated in the formula as $s^{1}_{t}=s^{2}_{t-1}$, $s^{2}_{t}=s^{3}_{t-1}$, $s^{3}_{t}=a_{t-1}$, which summarizes that the shelf-life of all products will decrease by one due to decaying. Then the $s_{t}$ will be used to compute the cost by equation (2). The second state transition occurs after the bakery interacts with the environment, returning daily demand information $d_{t}$:
@@
The goal of the inventory management problem is to find an optimal stationary policy (optimal $a_{t}$ given state $s_{t}$ at each period), such that the long-term expected cost $\mathbb{E}[\sum_{t=1}^{\infty}\gamma^{t-1}C_{t}]$ is minimized.
@@ III-B Implementation and Result Discussions
We first apply the adjusted Dyna-$Q$ algorithm (without transfer learning), with parameters $\alpha=0.3,\,\gamma=0.9,\epsilon_{0}=0.4,\,\epsilon_{\min}=0.1,\,\,N_{0}=100,\,N_{\min}=10,\,\partial_{\epsilon}=7500,\,\partial_{N}=5000$. For comparison, we construct the model $M(s,\,a)$ by two either BNNs or Multi-Layer Perceptrons (MLPs) (one for cost function and another for state transition probability). MLP is a neural network consisting of multiple layers of interconnected neurons. Its parameters are optimized by back-propagation, and the prediction error is minimized by gradient descent [24]. Our proposed algorithm is implemented across 500 training episodes, and the average cost per iteration has been recorded. The training processes are recorded in Figure 2, which shows the average cost occurs per iteration, where we observe a stable cost at the end of the training horizon. Thus, the trained optimal policies are also stable.
@@
Next, we implement the adjusted Dyna-$Q$ algorithm with transfer learning. We investigate the following two scenarios, training and testing, respectively. Here, the bakery manager is required to form an ordering policy at an early stage after the new product is released, without prior knowledge of its demand distribution. We choose the parameters $\alpha=0.1,\,\gamma=0.9,\,\partial_{\epsilon}=1000,\,\partial_{N}=1000$ and $\sigma^{2}=5$.
Scenario 1 (Training): In this scenario, we set: $\epsilon_{0}=0.4,\,\epsilon_{\min}=0,\,N_{0}=100,\,N_{\min}=0$. We first implement the method in Section II.$B$ with $h=10$ and incorporate transfer learning into the model of adjusted Dyna-$Q$. Then the training starts right at iteration $t=1$, when the new product, Boule 400g, has just been released and its demand is observed. The training horizons are set as one month i.e., $T=30$. The total cost of the training will be recorded, and its average value will be over 100 training episodes. We also record the variance of total cost over all episodes. Table II records the experiment results where the “Shortage percentage” records the proportion of days within a month when the unfulfilled demand occurs.
@@
Algorithm Transfer learning Average total cost Shortage percentage Daily average inventory holding Variance in total cost
@@
Algorithm Transfer learning Average total cost Shortage percentage Daily average inventory holding Variance in total cost
@@
From the experimental results of Scenario 1 recorded in Table II, we observe that incorporating transfer learning does not uniformly improve the algorithms’ performance. Regarding total inventory cost generated in the first month, the three algorithms do not show a distinct difference in Scenario 1, with or without transfer learning. Note that the basic Dyna-$Q$ is the most superrior, with the total cost during the training period at 61.86. However, incorporating transfer learning helps to stabilize all the algorithms through total cost variance reduction by at least $31.6\%$.
From the experimental results of Scenario 2 recorded in Table III, we observe that the adjusted Dyna-$Q$ algorithm with or without transfer learning performs better than all the other algorithms. In addition, incorporating transfer learning helps the adjusted Dyna-$Q$ algorithm perform the best on testing. Another observation is that the algorithms with transfer learning perform more stably with significantly lower variance in total cost on testing, than those without transfer learning. Finally, we observe the phenomenon that the adjusted Dyna-$Q$ algorithm produces policies reducing the shortage percentage by holding more inventory on average.
[…]
Figure 3: Transition probability comparison
@@ IV Conclusion
The paper presents a new RL algorithm for inventory management of the newly launched products where historical demand information is unavailable. The proposed adjusted Dyna-$Q$ algorithm is built upon the classic Dyna-$Q$ algorithm, aiming to strike a balance between model-based and model-free approaches during training. Additionally, based on the idea of transfer learning, a demand forecasting model is introduced, leveraging historical demand data from existing similar products to provide warm-start information for the adjusted Dyna-$Q$ algorithm.
A case study using real-life data is conducted to validate the proposed approach. The adjusted Dyna-$Q$ algorithm demonstrates significant benefits, including up to a 23.7% reduction in average daily cost compared to $Q$-learning and up to a 77.5% decrease in training time compared to classic Dyna-$Q$. Training and testing performances of incorporating generating transfer learning or not into the algorithm are shown and compared. The results show that the adjusted Dyna-$Q$ algorithm achieves the lowest total cost and relatively low shortage percentages compared to other algorithms. Algorithms that incorporate transfer learning exhibit greater stability and lower variance in total cost during testing.
For future research directions, several possibilities are suggested. First, further applications of the proposed approach in various inventory management contexts and industries could assess its versatility and scalability. Second, evaluating the adaptability of the approach to specific uncertainty patterns, such as seasonal demand fluctuations or sudden market changes, could provide insights into its robustness. Lastly, enhancing the demand prediction model to measure similarity between existing products and the new product could improve the incorporation of historical data from similar products with high similarity during training.